OpenStudy (anonymous):

Kyle has three short straws, four medium straws, and six long straws. If he randomly draws two straws, one at a time without replacement, what is the probability that both are short straws?

OpenStudy (anonymous):

how many total straws?

OpenStudy (anonymous):

That's what I'm trying to figure out myself.

OpenStudy (anonymous):

wait 13 straws in all right?

OpenStudy (anonymous):

very good.

OpenStudy (anonymous):

on the first draw, what is the prob of getting a short one?

OpenStudy (anonymous):

on the 2nd draw, remember that the number of short ones and the total number of straws have decreased by 1

OpenStudy (anonymous):

3.846%. The first time he draws a stick, the chance of it being short is 3/13, but the second time the chance is 2/12, because no replacement. so the probability is them multiplied, so 6/156=1/26=3.846%.

OpenStudy (anonymous):

on the first draw it is 3/13, which cannnot be reduced. on the 2nd draw it is 2/12=16 multiply the 2 togther and you get the prob of this outcome 3/13*(1/6)=3/78 or as wizguy said, 3.846%